OpenStudy (anonymous):

Three fair coins are tossed. What is the probability that at least one is a tail? (Enter the probability as a fraction.)

OpenStudy (anonymous):

1/3?

OpenStudy (anonymous):

If you flip a coin three times, how many different outcomes can happen? (i.e. HHH, TTT, HTH, etc,)

OpenStudy (dan815):

Prob of atleast 1, meaning 1 and 2 and 3 tails= 1-prob of no tails

OpenStudy (dan815):

Prob of all heads = (1/2)^3 Hence Prob of atleast 1 tail = 1-(1/2)^3 = 7/8

OpenStudy (anonymous):

ok thanks,